OpenStudy (anonymous):

The amount of power, P, generated by a windmill varies directly with the cube of the wind speed, w. A windmill on Oahu, Hawaii, produces 7300 kilowatts of power when the wind speed is 32 milers per hour. Answer as fractions or round to the nearest whole number. a. Express the power as a function of wind speed.

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (anonymous):

@ryan12345 thanks.. Can you help me with the problem?

OpenStudy (anonymous):

At first I was thinking that they just wanted w^3 = p but that doesnt jive with the sample data 32^3 != 7300 cube of 32 is 32768 so I'm not understanding the question I guess sorry

OpenStudy (anonymous):

@kantalope Thanks for trying dude. Yeah Im not sure either I neen to find K I think its something like L=k/x^2 But thats not working.

OpenStudy (anonymous):

If k is some constant p=(w^3)k p is still varying directly

OpenStudy (anonymous):

k is .22 or about a 1/4